please help find the correct solutions for question 1-5, i attempted to answer correctly but my teacher added this in the instructions

"To write these formulas, remember the following formula for classical probability:

?(?) = ? = ?????? ?? ???? ????? ? ??????/
? ????? ?????? ?? ???????? ???????? ?? ?h? ?????? ?????

Each of the questions involves a different “X”, and some of them involve different “T”s. Think through each question carefully to make sure you identify these values correctly."

and i couldnt figure out what value to have to be T, any help is appreciated!!!Household Income ($1000s) Education Level Not H.S. Graduate H.S. Graduate Some College Bachelors Degree Beyond Bach. Degree $0 -24.9 4,207 4,917 2,807 885 290 13,106 $25 -49.9$50-74.9$75 -99.9 1,389 5,027 4,678 2,848 1,274 15,216 3,459 6,850 5,258 2,094 829 18,490 539 2,637 3,250 2,581 1,241 10,248 $100+ 367 2,668 4,074 5,379 4,188 16,676 Total 9,961 22,099 20,067 13,787 7,822 73,736 Total What is the probability of a person not being a high school Question1 graduate? Answer 13.51% What is the probability of a person having a bachelors degree or Question 2 more education? Answer: 29.31% What is the probability that a person has a bachelors degree Question 3and makes more than $100,000 a year? Answer What is the probability of a person having an income belovw Question 4 S50,0002 Answer: 42.85% What is the probability that someone with some college or more Question 5 makes $75,000 or more? Answer 9.59%
Answer #1

Q1 The probability of a person not being a high school graduate is

p=Total person who is not high school graduate/ Total person

p= 9961/ 73736 = 0.1351

Q2 The probability of a person having a bachelor's degree or more education is

p=Total person whose education having a bachelor's degree or more education / Total person

p= (13787+7822)/ 73736 =0.2931

Q3 The probability of a person having a bachelor's degree and makes more than $100,000 a year is

p=Total person who taken bachelor's degree and makes more than $100,000 a year / Total person

p= 5379 / 73736 =0.0729

Q4 The probability of a person having an income below $50,000

p=Total person having an income below $50,000/ Total person

p= (13106+18490)/ 73736 =0.4285

Q5 The probability that someone with some college or more makes $75,000 or more

p=Total person study some college or more and makes $75,000 or more/ Total person

p=(10248+16676)/ 73736 =0.3651